7. A solution is made of two volatile solutes: Chemical A (with a pure vapor pressure of 80.0 mm Hg) and Chemical B (with a purePlease help is due in 30 mins, will rate, Thank you!
Answer #1

Given, for chemical A, PA0 = 80 mm Hg, number of moles of A, nA = 3

for chemical B, PB0 = 100 mm Hg, number of moles of B, nB  = ?

We know, Partial pressure of A, PA = XA x PA0 (XA = represents the mole fraction of component A)

=> PA = [3/(3+nB] x 80 mm Hg ----(i)

Partial pressure of B, PB = XB x PB0 (XB = represents the mole fraction of component B)

=> PB = [nB/(3+nB] x 100 mm Hg ----(i)

Now, Total pressure, PT = PA + PB

=> 98.8 mm Hg = [3/(3+nB] x 80 mm Hg + [nB/(3+nB] x 100 mm Hg

=> 98.8 = [240/(3+nB)] + [100nB/(3+nB)]

=> 98.8 x (3+nB) = 240 + 100nB

=> 296.4 + 98.8nB = 240 + 100nB

=> 1.2nB = 56.4

=> nB = 47 moles

Therefore 47 moles of B must be there.